At a candy store, there are 200 pieces of bubble gum in a bubble gum machine.
* Forty pieces are blue.
* Thirty-five pieces are red.
* Thirty pieces are green.
* Fifty pieces are yellow.
* Forty-five pieces are white.
If two pieces of bubble gum are purchased, what is the probability that both pieces will be yellow?

Answer by stanbon(75887)

At a candy store, there are 200 pieces of bubble gum in a bubble gum machine.
* Forty pieces are blue.
* Thirty-five pieces are red.
* Thirty pieces are green.
* Fifty pieces are yellow.
* Forty-five pieces are white.
If two pieces of bubble gum are purchased, what is the probability that both pieces will be yellow?
----
P(2 yellow) = 50C2/200C2 = 1225/19900 =49/796
============
Cheers,
Stan H.